undeadknight26 (undeadknight26):

Identify the sequence graphed below and the average rate of change from n = 1 to n = 3. coordinate plane showing the point 1, 8, point 2, 4, point 4, 1, and point 5, .5 an = 8(one half)n − 1; average rate of change is −3 an = 10(one half)n − 1; average rate of change is 3 an = 8(one half)n − 1; average rate of change is 3 an = 10(one half)n − 1; average rate of change is −3

OpenStudy (solomonzelman):

the x-coordinate is the number of the term, the y-coordinate is the term itself

undeadknight26 (undeadknight26):

Btw i think its D

OpenStudy (solomonzelman):

are you sure?

OpenStudy (solomonzelman):

sure that it is D?

undeadknight26 (undeadknight26):

I thought it was either A or D...Not 100% sure though

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle a_n=(a_1)\times {\rm r}^{n-1} }\)

undeadknight26 (undeadknight26):

Im quite sure its A now...

OpenStudy (solomonzelman):

the common ratio is 1/2 (you can see that every next y-coordinate is twice smaller than the previous one) and the first term is 8 (that is given explicitly by the first point) \(\large\color{black}{ \displaystyle a_n~~=~~(a_1)~~\times ~~{\rm r}~~^{n-1} }\) \(\large\color{black}{ \displaystyle a_n~~=~~~~8~~~~~\times ~~{\rm \frac{1}{2}}~~^{n-1} }\)

OpenStudy (solomonzelman):

yes, A is better:)

OpenStudy (solomonzelman):

you can tell the rate is negative (because it is an exponential decay) and you know the first term is 8

OpenStudy (solomonzelman):

and now know that r=1/2

undeadknight26 (undeadknight26):

That makes sense :D I actually graphed it using graphing technology!

undeadknight26 (undeadknight26):

Mind checking 2 more for me sire?

OpenStudy (solomonzelman):

sure, cool..... and the function you are dealing with comes out to be \(\large\color{black}{ \displaystyle y~~=8\times{\rm \left(\frac{1}{2}\right)}^{x-1} }\) \(\large\color{black}{ \displaystyle y~~=8\times{\rm \left(\frac{1}{2}\right)}^{x}\times \left(\frac{1}{2}\right)^{-1} }\) \(\large\color{black}{ \displaystyle y~~=8\times{\rm \left(\frac{1}{2}\right)}^{x}\times \left(2\right) }\) \(\large\color{black}{ \displaystyle y~~=16{\rm \left(\frac{1}{2}\right)}^{x}}\)

OpenStudy (solomonzelman):

this is just the function in terms of x.... I can check it, but I had a request to help with precalc, when I am done I will do my best to come back

undeadknight26 (undeadknight26):

Oh that makes sense...